The same exam was given to all the students in the sixth grade, but they were...

  1. The same exam was given to all the students in the sixth grade, but they were divided into 4 classrooms. Each classroom had a different distraction taking place except for Room 1 that was the control group and free of noise. Using the scores ranging from 0 to 60 of the students below, state a conclusion about the nature of testing practices used.

Control

60 degrees

Classical Music

Next to Cafeteria

31

53

27

29

16

39

36

18

11

41

32

16

19

39

14

20

43

54

40

7

49

41

60

12

4

9

47

20

11

17

2

45

34

13

27

33

39

8

5

53

60

32

55

23

40

29

53

24

57

35

27

18

59

45

4

49

43

42

50

29
Answer #1

The table below gives the data which has been calculated from the given figures

Total 516 497 479 396
n 15 15 15 15
Mean 34.40 33.13 31.93 26.40
Sum Of Squares 4791.6 3143.7 6179.2 2858.6
Variance 342.26 224.55 441.37 204.19
Standard deviation 18.50 14.98 21.01 14.29

The Hypothesis:

H0: There is no difference between the mean scores due to the 4 different testing methods.

Ha: There is a difference between the mean scores due to the 4 different testing methods.

The ANOVA table is as below. the p value is calculated for F = 0.62 for df1 = 4 and df2 = 56

The Fcritical is calculated at = 0.05 (default) for df1 = 4 and df2 = 56

Source SS DF Mean Square F Fcv p
Between 558.86 3.00 186.29 0.62 2.77 0.6082
Within/Error 16973.18 56.00 303.09
Total 17532.04 59.00

The Decision Rule:

If Ftest is > F critical, Then Reject H0.

Also if p-value is < , Then reject H0.

The Decision:

Since Ftest (0.62) is < F critical (2.77), We Fail to Reject H0.

Also since p-value (0.6082) is > (0.05), We reject H0.

The Conclusion: There isn't sufficient sufficient evidence at the 95% level of significance to conclude that there is a difference between the mean scores due to the 4 different testing methods.

___________________________________________________________________

Calculations For the ANOVA Table:

Overall Mean = (516 + 497 + 479 + 396) / 60 = 31.47

SS treatment = SUM n* ( - overall mean)2 = 15 * (34.4 - 31.47)2 + 15 * (31.93 - 31.47)2 + 15 * (33.13 - 31.47)2 + 15* (26.4 - 31.47)2 = 558.86

df1 = k - 1 = 3 - 1 = 3

MSTR = SS treatment/df1 = 558.86 / 3 = 186.29

SSerror = SUM (Sum of Squares) = 4791.6 + 3143.7 + 6179.2 + 2858.6 = 16973.18

df2 = N - k = 60 - 4 = 56

Therefore MS error = SSerror/df2 = 16973.18 / 56 = 303.09

F = MSTR/MSE = 186.29 / 303.09 = 0.62

____________________________________________________